OpenStudy (dls):

Find the approximate value of cube root of 127.

OpenStudy (dls):

Attempt: \[\Huge f(x)=(x)^\frac{1}{3}\] \[\LARGE f(x+\Delta x)=f(x)+\Delta x(f'(x))\] this is the derived formula.. substituting x=125 and delta x=2 \[\LARGE f(127)=f(125)+2 \times \frac{1}{3} \times (\frac{1}{125})^{\frac{2}{3}}\]

OpenStudy (dls):

\[\LARGE \frac{50}{3}+5=>\frac{65}{3}?\]

OpenStudy (dls):

hmm seems wrong :P

OpenStudy (anonymous):

@DLS \(2\times\frac13\times(\frac1{125})^\frac23\ne\frac{50}3\)

OpenStudy (dls):

2/3 x 1/25 sorry!!

OpenStudy (dls):

2/75

OpenStudy (anonymous):

You incorrectly simplified \((1/125)^{2/3}=25\) rather than \(1/25\)

OpenStudy (dls):

\[\LARGE 5+\frac{2}{75}=377/75\]